CNN wants to determine the proportion of the population in california that takes social distancing seriously. So they wants to have a 90% confidence interval whose length is 0.1. They pollster charges $10 dollars per respondent. How much money must they editor budget for the poll?
Answer #1

Ans:

Confidence interval length=0.1

Margin of error=0.1/2=0.05

Minimum sample size required,n=1.645^2*0.5*(1-0.5)/0.05^2=270.60

So,budget required=10*271=2710